OpenStudy (anonymous):

The function f is one to one. Find its inverse. State the domain of f and find its range using f^-1 f(x)= -9x/x-1 f^-1(x)=_______ domain f=______ range f=_______

OpenStudy (anonymous):

x-1=0 x = 1 plug x = 1 find y -9(1)/1-0 =0 \[x(y-1)=-9y\] y =0\[y=\frac{-9x}{x-1}\]Substitution x=y\[x=\frac{-9y}{y-1}\]\[x(y-1)=-9y\] \[xy-x=-9y\]\[xy+9y=x\]\[y(x+9)=x\]\[y=\frac{x}{x+9}\] \[f^-1=\frac{x}{x+9}\]domain f=1 range f=0